OpenStudy (anonymous):

Medal award! A bicycle with 19 in diameter wheels has its gear set so that the chain has a 5-in. radius on the front sprocket and 4-in radius on the rear sprocket. The cyclist pedals at 105rpm.

OpenStudy (anonymous):

Hey can u help me with some math questions please

OpenStudy (anonymous):

bella280, sure which question?

OpenStudy (anonymous):

c for front sprocket=2 pi r=2pi*5=10 pi no. of revolutions per minute=105 distance covered by front sprocket (circular)=105*10pi=1050 pi c for rear sprocket =2 pi r=2*pi*4=8 pi no. of revolutions by rear sprocket per minute=1050 pi/8pi=525/4 diameter of the wheel=19in radius=19/2=9.5 in c for wheel (distance in 1 revolution)=2 pi r=2 pi*9.5=19 pi distance in 525/4 revolutions=19*525/4=?

OpenStudy (anonymous):

so 2493.75in the linear speed?

OpenStudy (anonymous):

This is linear speed 2493.75 in / minute

OpenStudy (anonymous):

for some reason it says it's wrong